OpenStudy (clamin):

PLEASE HELP!!! MEDAL!! Below are five number sequences. Which ones are geometric sequences? a. 1/3, 2/9, 3/274, 4/81, 5/343.... b. 3, -15, -33, -51, -69... c.3, -12, 48, -192, 768..... d. 1, 1/2, 1/4, 1/8, 1/16.... e. 2, 3, 5, 9, 17...

OpenStudy (clamin):

i made a mistake on letter a.. its supposed to be 3/27 not 3/274

OpenStudy (clamin):

I think there 1 or more answers

OpenStudy (studygurl14):

Divide each number in the sequence by the number before it. If the numbers in the same list result in the same thing, it is a geometric sequence.

OpenStudy (clamin):

I dont get it

OpenStudy (clamin):

I dont get it

OpenStudy (studygurl14):

For example. The first sequence: 1/3, 2/9, 3/27, 4/81, 5/343 2/9 divided by 1/3 = 2/3 3/27 divided by 2/9 = 1/2 2/3 \(\neq\) 1/2, so that is not a geometric sequence. Understand the process now?

OpenStudy (clamin):

Yes thank you @studygurl14